# Collision Energy Dependence of Particle Ratios and Freeze-out Parameters in Ultra Relativistic Nucleus Nucleus Collisions

### v1

@@ III.3 ENERGY DEPENDENCE OF THERMAL FREEZE-OUT PARAMETERS ($T$, $\mu_{B}$, $\mu_{s}$)
In order to highlight the energy dependence of the freeze-out parameters of the model, we have shown in Figs. 17 and 18 the $\sqrt{s_{NN}}$ dependence of $T$, $\mu_{B}$, $\mu_{s}$, and $\mu_{Q}$. The energy dependence of $\xi_{s}$ has already been discussed in the previous section shown in Fig 7. Fig. 17 shows how the system’s temperature varies with collision energy in URNNC. The two separate curves are for baryonic and mesonic degrees of freedom as they are found to freeze-out at different stages of the system’s evolution as discussed earlier also. However, both curves saturate beyond $\sqrt{s_{NN}}\sim 60\,\text{GeV}$ at 172 MeV and 162 MeV, respectively. This may indicate the onset of nuclear transparency around this energy, with lesser stopping of the colliding nuclear medium and a large production of secondary hadrons during the evolution of the system Bashir2015 . In Fig. 18, we have shown the energy dependence of $\mu_{B}$ and $\mu_{s}$ extracted from the thermal fits of particle ratios. The $\mu_{B}$ is seen to initially decrease rapidly with $\sqrt{s_{NN}}$ and later approach a negligibly small value beyond 200 GeV. The $\mu_{s}$, which is obtained by applying the strict strangeness conservation criteria in the hot and dense system formed in URNNC, initially shows a slight increasing trend up to 10 GeV, reaching a value of approximately 50 MeV, and then shows a decreasing trend to approach almost zero value beyond 200 GeV. Both $\mu_{B}$ and $\mu_{s}$ thus approach zero as the matter becomes more and more baryon symmetric with increasing $\sqrt{s_{NN}}$, which again is an indicator of the nuclear transparency effect setting in at the higher RHIC energies and beyond up to LHC energies.
@@
Fig. 19 illustrates a smooth parabolic dependence of the chemical freeze-out stages in the $T-\mu_{B}$ plane. The vertical axis represents the chemical freeze-out temperature, while the horizontal axis shows the baryon chemical potential, both obtained using our model calculations for different $\sqrt{s_{NN}}$. The red and blue curves in the figure again correspond to the baryonic and mesonic degrees of freedom, respectively. At RHIC energies, the chemical freeze-out temperatures $T_{che}\left(\approx 163-170\,\text{MeV}\right)$ are close to the hadron-QGP phase transition temperature predicted by lattice gauge theory Bashir2015 ; Baban2013 ; Tiwari2013 thus indicating that freeze-out occurs almost immediately after the phase transition.
@@
Our model results align closely with the findings from some previous studies on freeze-out parameters in heavy-ion collisions Cleymans2006C ; Munzinger2001 ; Adams2005E ; Baran2003 ; Florkowski2001fp ; Chatterjee2015 ; Alba2020ChemicalFP . The freeze-out temperatures ($T$) and baryon chemical potentials ($\mu_{B}$) for RHIC energies have been earlier reported, showing a range of $T$ values from 163 MeV to 177 MeV and $\mu_{B}$ values from 24 MeV to 29 MeV for $\sqrt{s_{NN}}=200\,\text{A GeV}$, and $T$ values from 165 MeV to 174 MeV and $\mu_{B}$ values from 38.1 MeV to 46 MeV for $\sqrt{s_{NN}}=130\,\text{A GeV}$ Munzinger2001 ; Cleymans2005C .
Figure 17:
Figure 18:
Figs. 17-18 Variation of $T$ and $\mu_{B}$, $\mu_{s}$ with $\sqrt{s_{NN}}$
Figure 19: Variation in the $T-\mu_{B}$ plane
IV SUMMARY AND CONCLUSION
@@
In summary we have systematically applied the statistical VDW-HRG model in the framework of grand canonical ensemble to analyze relative particle yields in ultra-relativistic nucleus-nucleus collisions. We have used the values of the model parameters representing the attractive and repulsive parts of interaction, which are obtained from the ground state properties of nuclear matter. Using an established ansatz form for the chemical freeze-out temperatures and baryon chemical potential, we have determined the thermo-chemical freeze-out parameters from the $\bar{p}/p$ and $K^{-}/K^{+}$ ratios. We have found that there is a need to consider separate freeze-out temperatures for baryons and mesons (Kaons and pions). This is justified by their different interaction cross-sections in the system. We have further explained strange particle-to-pion ratios successfully, including the experimentally observed hump in $K^{+}/\pi^{+}$, by introducing a strangeness imbalance factor ($\xi_{s}$) derived from $K^{-}/\pi^{-}$ ratios. Using the model, we have quite successfully explained the experimental results for several like and unlike mass hadron ratios over a wide range of collision energies. The experimentally observed correlation between $K^{-}/K^{+}$ and $\bar{p}/p$ ratios is well explained by our model calculations. A common feature of all ratios is that they tend to saturate at large collision energies. This is a clear indication of a nearly constant temperature and chemical potential (though vanishingly small) at higher collision energies, i.e., beyond 200 GeV. In the case of unlike mass particle ratios, the initially increasing temperature with $\sqrt{s_{NN}}$ favors a relatively rapid increase in the production of heavier particles. The decreasing chemical potential also favors the production of antibaryons, while for baryons, it does the opposite. We have illustrated the energy dependence of thermal freeze-out parameters $T$, $\mu_{B}$, and $\xi_{s}$, indicating the complex nature of particle production at high energies.

### v2

@@ IV.3 ENERGY DEPENDENCE OF THERMAL FREEZE-OUT PARAMETERS ($T$, $\mu_{B}$, $\mu_{s}$, $\mu_{Q}$)
In order to highlight the energy dependence of the freeze-out parameters of the model, we have shown in Fig. 9 the $\sqrt{s_{NN}}$ dependence of $T$, $\mu_{B}$, $\mu_{s}$, and $\mu_{Q}$ as extracted from the thermal fits of particle ratios. The energy dependence of $\xi_{s}$ has already been discussed in the previous section and shown in Fig. 4. The temperature curve saturates beyond $\sqrt{s_{NN}}\sim 50\,\text{GeV}$ at 164 MeV. As discussed before this may indicate the onset of nuclear transparency around this energy, where there is lesser stopping of the colliding nuclear medium and the bulk production of secondary hadrons takes place during the evolution of the system between the receding nuclei Bashir2015 . The $\mu_{B}$ is seen to initially decrease rapidly with $\sqrt{s_{NN}}$ and later approach a negligibly small value beyond 200 GeV. The $\mu_{s}$, which is obtained by applying the strict strangeness conservation criteria in the hot and dense system formed in URNNC, shows a decreasing trend from an initial value of $\sim$ 90 MeV to almost zero value beyond 200 GeV. Both $\mu_{B}$ and $\mu_{s}$ thus approach zero as the matter becomes more and more baryon symmetric with increasing $\sqrt{s_{NN}}$, which again is an indicator of the nuclear transparency effect setting in at the higher RHIC energies and further beyond up to LHC energies. In the same figure we have also shown the variation of the electric chemical potential $\mu_{Q}$. We find that there is an increasing trend from a very small initial value ($\sim$ -16 MeV) to almost zero towards higher energies as the system tends to become (almost) baryon symmetric.
Figure 9: Collision energy dependence of $\mu_{B}$, $\mu_{s}$, $\mu_{Q}$ and $T_{che}$.
@@
Fig. 10 illustrates a somewhat smooth parabolic dependence of the chemical freeze-out stages in the $T-\mu_{B}$ plane. The vertical axis represents the chemical freeze-out temperature, while the horizontal axis shows the corresponding baryon chemical potential. At RHIC energies, the chemical freeze-out temperatures $T_{\text{che}}$ ($\sim$ 160 - 164 MeV) are close to the hadron-QGP phase transition temperature predicted by lattice gauge theory Tiwari2013 ; Lysenko2024Chemical thus indicating that freeze-out occurs almost immediately after the phase transition. For the sake of comparison, on the same graph are also shown some earlier results. The point-like case results of Cleymans et. al. Cleymans2006C and Andronic et. al. Andronic2006 are shown along with the finite size case result of Poberezhnyuk et al. Poberezhnyuk2019Chemical who have used the baryonic hard-core radius value as 0.59 fm. The results of Cleymans et. al. are in close agreement with our results while the results of Andronic et. al. and Poberezhnyuk et. al. show some deviations at small and intermediate collision energies.
Figure 10: Chemical freeze-out line from different calculations
It may be noted that the extracted values of the freeze-out parameters in our case is dependent on the various particle ratios as obtained at different collision energies. On the other hand Poberezhnyuk et. al. have used the data on mean hadron multiplicities in various experiments to extract parameter values.
@@
Our current model results nevertheless, in general, align somewhat closely with the findings of some other previous studies. The freeze-out temperatures ($T$) and baryon chemical potentials ($\mu_{B}$) for RHIC energies have been earlier reported, showing a range of $T$ values from 151 MeV to 167 MeV and $\mu_{B}$ values from 50 MeV to 20 MeV for higher RHIC energies, $\sqrt{s_{NN}}=130-200\,\text{GeV}$ Andronic2017Decoding ; Cleymans2006C ; Munzinger2001 ; Adams2005E ; Baran2003 ; Florkowski2001fp ; Chatterjee2015 ; Alba2020ChemicalFP ; Lysenko2024Chemical ; Poberezhnyuk2019Chemical ; Cleymans2005C ; Andronic2006 ; Florkowski2001 ; Tawfik2013 ; Tawfik2014 .
V SUMMARY AND CONCLUSION
@@
In summary we have systematically applied the statistical VDW-HRG model in the framework of grand canonical ensemble to simultaneously describe several particle ratios in ultra-relativistic nucleus-nucleus collisions over a broad range of collision energy using a single set of model parameters. We have used a mean value of the parameter representing the attractive part of interaction obtained from the ground state properties of nuclear matter. The baryonic (antibaryonic) hard-core radius is treated as a free parameter. We have used an established ansatz form for the chemical freeze-out temperatures and baryon chemical potential. We have determined values of all freeze-out parameters from the (simultaneous) best theoretical fits obtained for the like-mass antibaryon to baryon ratio data sets using minimum chi squared fit method. We have explained the $K^{-}/K^{+}$, $\pi^{-}/\pi^{+}$ and proton to pion ratios quite well. Several other unlike-mass particle ratios (strange to non-strange) have also been well describes by the present calculation by introducing a strangeness imbalance factor ($\xi_{s}$). The experimentally observed correlation between $K^{-}/K^{+}$ and $\bar{p}/p$ ratios is well explained by our model calculations. A common feature of all ratios is that they tend to saturate at large collision energies. This is a clear indication of a nearly constant temperature and chemical potential (though vanishingly small) at higher collision energies, i.e., beyond 200 GeV. The decreasing chemical potential favours the production of antibaryons over baryons. We have illustrated the energy dependence of the freeze-out line and compared it with the previous results.